THE E ^{2}Π - X ^{2}Σ^{+} STATE OF CaH/D

\underline{K. TERESZCHUK}, I. GORDON and P.F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario Canada N2L 3G1.

Our group has recently investigated the E^{2}Π-X^{2} Σ^{+} band system of CaH/D, while searching for the elusive double minima electronic states of the D^{2} Σ^{+}-X^{2} Σ^{+} system. Emission spectra of CaH and CaD, in the visible region, have been obtained by means of a new discharge-furnace source. Calcium metal was placed into an evacuated, alumina tube and heated to a temperature of 600 ^{o}C in a tube furnace. A mixture of argon and hydrogen gas was allowed to flow through the tube and was then excited by a DC electrical discharge. The spectra were recorded using a photomultiplier tube with our Bruker IFS 120 HR spectrometer. We will report improved values for the vibrational and rotational constants of CaH/D.